# Blind Inverse Problem Solving Made Easy by Text-to-Image Latent Diffusion

### v1

@@ 4.3 Blind JPEG Decompression
We further compare our method with baselines on JPEG decompression, a particularly challenging task due to the non-linear, non-differentiable nature of JPEG compression. Unlike traditional settings, our experiments provide no task-specific knowledge, such as the compression algorithm, quantization table, or factors – algorithms rely solely on the measurement images.
Setup
We generate measurements using JPEG compression with a quantization factor of 2. We use a neural network with a 3-layer U-net [32] to parametrize the operator. The operator is initialized using Algorithm 2 with $M=8$ and $N=4$. We use phrases including “pixelated, low quality, jpeg artifacts, deformed, bad proportions, …” as negative prompt, whose full version is provided in the appendix.
Results
Table 3: Quantitative results on blind JPEG decompression task.
@@ Results
SelfDeblur 0.676 8.84 0.1959 0.703 8.89 0.1173
MPRNet 0.785 5.99 0.8008 0.769 5.93 0.6809
DeblurGANv2 0.473 21.48 0.2207 0.502 21.76 0.1527
BlindDPS 0.431 21.55 0.1791 0.397 20.87 0.2108
GibbsDDRM 0.841 13.91 0.2915 0.775 14.59 0.2915
LADiBI (Ours) 0.268 21.40 0.0172 0.315 21.12 0.0216
@@
4.4 Ablation Studies
Table 4: Ablation study on our proposed algorithm.
Removed Component LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
Text prompts 0.500 18.88 0.0483
MPGD guidance 0.470 17.80 0.0434
$\hat{\phi}$ Initialization 0.355 19.74 0.0250
LADiBI 0.315 21.12 0.0216
To showcase the effectiveness of each part of our algorithm, we conduct ablation study on AFHQ dataset with JPEG decompression task. In particular, we test the importance of suitable ad-hoc prompts, the use of MPGD guidance, and the operator initialization for neural network ${\mathcal{A}}_{\hat{\phi}}$. in-between diffusion steps. We keep SDEdit in all settings to encode the measurement information for fair comparisons. As shown in Table 4, each of the aforementioned components plays a significant role in our scheme and is indispensable for producing high quality results.
@@ 4.5 Style Preserving Deblurring
Figure 4: Qualitative results on blind deblurring Monet paintings.
In order to demonstrate our method’s ability to solve blind inverse problems from a wider range of image distributions, we present indicative examples of Gaussian and motion deblurring on Monet paintings in Figure 4. We use the same negative prompts as in the previous experiments and “a portrait of a person as a Monet style painting” as positive prompts. We notice that, although the images portray human faces, the baseline method using models trained on realistic human face images cannot accurately solve the problem, while our method effectively generates images consistent with both the measurement image and the painting style present in the ground truth image.
@@ 5 Conclusion
In this work, we propose LADiBI, a new training-free algorithm to solving blind inverse problems in image restoration using a large-scale pre-trained text-to-image diffusion model. Our method leverages text prompts as well as posterior guidance on intermediate diffusion steps to guide the generation process towards clean reconstructions based on degraded measurements, in cases where the degradation operation is unknown. Experimental results demonstrate that LADiBI achieves competitive performance on tasks with diverse operator and is also capable of covering a wider range of image distributions compared to baselines.
@@ Appendix A Additional Implementation Details
We offer more detailed information regarding the implementation setup for Algorithms 1 and 2. All experiments are conducted using an NVIDIA A6000 GPU. In an effort to make our methodology generalizable and extensible to other inverse problems, we attempt to maintain the same hyper-parameter values across tasks wherever possible. Each hyper-parameter value has been chosen after conducting preliminary experiments for a specific range and opting for the value that offers the best performance. Indicative examples of such experiments are shown in C. In particular, we run Algorithm 1 for $T_{s}=150$ timesteps while performing 4 repetitions as part of the time-traveling strategy. We encode the measurement as ${\bm{x}}_{T_{s}}$ by applying the forward diffusion process up to timestep 800. In parallel with the reverse diffusion process, we update $\hat{\phi}$ every 5 timesteps, each time conducting $K=150$ gradient steps. We adust the $c_{t}$ and the $\lambda_{\phi}$ parameters of Equations 8 and 9 to 30 and 2 respectively. In terms of the operator initialization process described by Algorithm 2, we make use of a batch of $N=4$ samples and run $M=8$ iterations, each comprising $T_{j}=60$ timesteps.
Figure 5: A schematic overview of Algorithm 1.
Figure 6: A schematic overview of Algorithm 2.
Additionally, schematic overviews of Algorithms 1 and 2 are presented in Figures 5 and 6.
Furthermore, there are some parameters in our approach for which employing a task-aware setup strategy is essential for state-of-the-art performance. These are the prompt sentences as well as the architecture of $\hat{{\phi}}$. Here we provide details with respect to these parameters according to each restoration task:
@@
  • Positive prompts: “a clear headshot of a person/animal”
  • Negative prompts: “shaken image, motion blur, pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
@@
  • Positive prompts: “a clear headshot of a person/animal”
  • Negative prompts: “blurry, gaussian blur, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
@@
  • Positive prompts: “a clear headshot of a person/animal”
  • Negative prompts: “pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, jpeg artifacts, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions”
  • $\hat{\phi}$ architecture: A neural network with a typical 3-layer U-net [32] architecture. Each layer consists of 2 convolutional blocks of size $3\times 3$ with ReLU activations and number of input and output channels ranging from 32 to 128.
@@ Appendix B Additional Results
Table 5: Additional quantitative results on blind JPEG decompression task.
Method FFHQ AFHQ
@@
Pan-l0 0.787 12.72 0.2190 0.825 13.15 0.2168
Pan-DCP 0.710 14.91 0.2318 0.673 18.27 0.2472
SelfDeblur 0.676 8.84 0.1959 0.703 8.89 0.1173
MPRNet 0.785 5.99 0.8008 0.769 5.93 0.6809
DeblurGANv2 0.473 21.48 0.2207 0.502 21.76 0.1527
BlindDPS 0.431 21.55 0.1791 0.397 20.87 0.2108
GibbsDDRM 0.841 13.91 0.2915 0.775 14.59 0.2915
LADiBI (Ours) 0.268 21.40 0.0172 0.315 21.12 0.0216
@@
We present more detailed experimental results on all benchmark tasks. In specific, Table 5 shows quantitative results on the JPEG decompression task, including the optimization-based Pan-l0 and Pan-DCP methods. Moreover, Figures 7, 8 and 9 offer more qualitative comparisons against baseline methods on motion deblurring, gaussian deblurring and JPEG decompression respectively. Finally, in Fig. 10 we include comparisons against selected baseline methods on the 3 benchmark tasks using a set of Monet and Van Gogh portrait paintings as ground truth images.
@@ Appendix C Additional Ablation Study
We carefully adjust each parameter of our methodology to the value that offers the best overall results by conducting comparative experiments and in this section we offer some indicative examples of such tests. All hyper-parameter tuning experiments have been performed on the motion deblurring task using the AFHQ dataset.
Figure 11 presents both the LPIPS and KID score for the value of the timestep at which we encode the measurement using the forward diffusion process. Both metrics form a U-shape curve, which results from the following fact about encoding the measurement image; if we make the encoding timestep too large, most of the information about the measurement has been replaced by noise which does not allow the sampling process to leverage useful features of the degraded image. On the other hand, if the timestep is too small, the sampling process is not equipped with enough scheduled variance to reach the target prior distribution.
@@
With respect to ablation studies, we are interested in evaluating the performance of our algorithm when using a task-agnostic implementation for ${\mathcal{A}}_{\hat{\phi}}$. Hence, we tested the neural network architecture employed for non-linear tasks at linear inverse problems and we showcase qualitative results in Figure 13. We observe that, although performing worse than LADiBI employed with a linear operator architecture, the task-agnostic method is capable of producing estimates of decent quality while also preserving a design that allows applicability to any image restoration task.
Appendix D Limitations and Future Works

### v2

@@ 4.3 Blind JPEG Decompression
We further compare all methods on JPEG decompression, a particularly challenging task due to its non-linear, non-differentiable nature. Unlike traditional settings, our experiments provide no task-specific knowledge, such as the compression algorithm, quantization table, or factors – algorithms rely solely on the measurement images.
We generate measurements using JPEG compression with a quantization factor of 2. We use a neural network with a 3-layer U-net (Ronneberger et al., 2015) to parametrize the operator. The operator is initialized using Algorithm 2 with $M=8$ and $N=4$.
Table 3: Quantitative results (the mean and standard deviation) on blind JPEG decompression task.
@@
Pan-I0 0.787 $\pm$ 0.084 12.72 $\pm$ 2.53 0.2189 0.825 $\pm$ 0.083 13.15 $\pm$ 2.68 0.2168
Pan-DCP 0.710 $\pm$ 0.067 14.91 $\pm$ 2.63 0.2318 0.673 $\pm$ 0.055 18.27 $\pm$ 1.60 0.2472
SelfDeblur 0.676 $\pm$ 0.054 8.84 $\pm$ 1.93 0.1959 0.703 $\pm$ 0.052 8.89 $\pm$ 1.72 0.1173
MPRNet 0.785 $\pm$ 0.048 5.99 $\pm$ 1.81 0.8008 0.769 $\pm$ 0.048 5.93 $\pm$ 1.78 0.6809
DeblurGAN 0.473 $\pm$ 0.058 21.48 $\pm$ 1.85 0.2207 0.502 $\pm$ 0.061 21.76 $\pm$ 1.84 0.1527
BlindDPS 0.431 $\pm$ 0.079 21.55 $\pm$ 1.96 0.1791 0.397 $\pm$ 0.069 20.87 $\pm$ 1.89 0.2108
BIRD 0.406 $\pm$ 0.047 20.68 $\pm$ 1.09 0.0525 0.425 $\pm$ 0.068 21.08 $\pm$ 1.50 0.0673
GibbsDDRM 0.841 $\pm$ 0.057 13.91 $\pm$ 1.26 0.2915 0.775 $\pm$ 0.059 14.59 $\pm$ 1.48 0.2915
Ours 0.268 $\pm$ 0.070 21.40 $\pm$ 1.18 0.0172 0.315 $\pm$ 0.075 21.12 $\pm$ 1.27 0.0216
@@
4.4 Ablation Study
Table 4: Ablation study on JPEG decompression.
Ablation LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
W/o text prompts 0.508 19.30 0.0319
W/o negative prompts 0.440 19.44 0.0242
Using generic $\bm{c}$$_{-}$ 0.425 19.48 0.0255
Using task-irrelevant $\bm{c}$$_{-}$ 0.406 19.40 0.0237
W/o MPGD guidance 0.403 19.44 0.0274
W/o regularization 0.307 21.35 0.0172
W/o $\hat{\phi}$ Initialization 0.289 20.17 0.0170
LADiBI 0.262 21.20 0.0132
Figure 5: Qualitative results on blind deblurring Monet paintings.
To showcase the effectiveness of each part of our algorithm, we conduct ablation study on AFHQ dataset with JPEG decompression task. In particular, we test the importance of suitable ad-hoc prompts, the use of MPGD guidance, the MPGD regularization term, and the operator initialization for neural network $^{\mathcal{A}}_{\hat{\phi}}$. We keep SDEdit in all settings to encode the measurement information for fair comparisons. As shown in Table 4, each of the aforementioned components plays a significant role in our scheme and is indispensable for producing high quality results.
@@ 4.5 Style Preserving Deblurring
In order to show our method’s applicability on a wider range of image and operator distributions, we present indicative examples of Gaussian and motion deblurring on Monet paintings in Figure 5. Additional operators and image distributions demonstrations are available in Appendix B. We use the same negative prompts as in the previous experiments and “a portrait of a person as a Monet style painting" as positive prompts. We notice that, although the images portray human faces, the baseline method using models trained on realistic human face images cannot accurately solve the problem, while our method effectively generates images consistent with both the measurement image and the painting style present in the ground truth image.
@@ 5 Conclusion
In this work, we propose LADiBI, a new training-free algorithm to solving blind inverse problems in image restoration using large-scale pre-trained text-to-image diffusion models. With unknown degradation operators, our method leverages text prompts as well as posterior guidance on intermediate diffusion steps to restore desired images based on the measurements. Experiments demonstrate that LADiBI’s effectiveness on diverse operator and image distributions.
@@ Appendix A Additional Implementation Details
In this section, we offer more detailed information regarding the implementation setup for Algorithms 1 and 2. All experiments are conducted using an NVIDIA A6000 GPU. In an effort to make our methodology generalizable and extensible to other inverse problems, we attempt to maintain the same hyper-parameter values across tasks wherever possible. Each hyper-parameter value has been chosen after conducting preliminary experiments for a specific range and opting for the value that offers the best performance. Indicative examples of such experiments are shown in C.
In particular, we run Algorithm 1 for $T_{s}=150$ timesteps while performing 4 repetitions as part of the time-traveling strategy. We encode the measurement as ${\bm{x}}_{T_{s}}$ by applying the forward diffusion process up to timestep 800. In parallel with the reverse diffusion process, we update $\hat{\phi}$ every 5 timesteps, each time conducting $K=150$ gradient steps. We adjust the $c_{t}$ and the $\lambda_{\phi}$ parameters of Equations 8 and 9 to 30 and 2 respectively. In terms of the operator initialization process described by Algorithm 2, we make use of a batch of $N=4$ samples and run $M=8$ iterations, each comprising $T_{j}=60$ timesteps.
We also take into consideration that the targets and measurements reside in the $256\times 256$ pixel space, whereas Stable Diffusion v1 operates on images of pixel size $512\times 512$. To address this disparity, we initially upsample the measurement using bilinear interpolation in order to transform it to a $512\times 512$ image, and then downsample the resulting $x_{0}$ to map the final estimate back to the original image space.
The schematic overviews of Algorithms 1 and 2 are presented in Figures 2 and 6.
In addition, there are some parameters in our approach for which employing a task-aware setup strategy is essential for state-of-the-art performance. These are the prompt sentences as well as the architecture of $\hat{{\phi}}$. Here we provide details with respect to these parameters according to each restoration task:
@@
  • Positive prompts: “a clear headshot of a person/animal"
  • Negative prompts: “shaken image, motion blur, pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions"
@@
  • Positive prompts: “a clear headshot of a person/animal"
  • Negative prompts: “blurry, gaussian blur, lowres, text, error, cropped, worst quality, blurry ears, low quality, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions"
@@
  • Positive prompts: “a clear headshot of a person/animal"
  • Negative prompts: “pixelated, lowres, text, error, cropped, worst quality, blurry ears, low quality, jpeg artifacts, ugly, duplicate, morbid, mutilated, text in image, DSLR effect, poorly drawn face, mutation, deformed, blurry, dehydrated, blurry hair, bad anatomy, bad proportions, disfigured, gross proportions"
  • $\hat{\phi}$ architecture: A neural network with a typical 3-layer U-net [Ronneberger et al., 2015] architecture. Each layer consists of 2 convolutional blocks of size $3\times 3$ with ReLU activations and number of input and output channels ranging from 32 to 128.
We conduct quantitative experiments on FFHQ $256\times 256$ [Karras et al., 2019] and AFHQ $256\times 256$ [Choi et al., 2020]. Images in FFHQ are publicized under Creative Commons BY 2.0, Creative Commons BY-NC 2.0, Public Domain Mark 1.0, Public Domain CC0 1.0, or U.S. Government Works license and AFHQ is publicized under Attribution-NonCommercial 4.0 International license.
Figure 6: A schematic overview of general operator initialization (Algorithm 2).
@@ Appendix B Additional Results
In this section, we present additional experimental results on all benchmarks as well as additional tasks and data distributions. Table 5 include the standard deviation of the linear benchmark results. Figures 7, 8 and 9 offer more qualitative comparisons against baseline methods on motion deblurring, gaussian deblurring and JPEG decompression respectively. In Figure 10, we include comparisons against selected baseline methods on the 3 benchmark tasks using a set of Monet and Van Gogh portrait paintings as ground truth images.
We also show additional demonstrations of applying LADiBI to solve colorization and non-linear deblurring problems. Figures 12 and 12 substantiate our claim that our algorithmic scheme is flexible enough to adapt to various inverse tasks. Additionally, thanks to the broad range of image distributions encapsulated by the pretrained model, as well as the minimal assumptions imposed by the proposed methodology, our approach is able to accurately solve inverse problems for a broad range of images. Figures 14 and 14 demonstrate our method’s capability to perform Gaussian deblurring, motion deblurring as well as JPEG decompression for images depicting cars and landscapes.
Finally, to assess the real-world performance of our method, we test it on samples from the RealBlur dataset [Rim et al., 2020], which contains naturally captured blurry images using GoPro in real life settings. This dataset presents realistic blur patterns that are significantly more complex and diverse compared to the synthetically generated benchmarks. Figure 15 presents qualitative demos, which highlight our method’s ability to recover sharp structures and textures under various real-world blur conditions. These results demonstrate the applicability of our approach to real inverse problem solving tasks and support its generalization capability.
Table 5: Quantitative results on motion and Gaussian deblurring on FFHQ dataset. Mean $\pm$ standard deviation shown.
Method Motion Gaussian
@@
DPS w/ GT kernel 0.164 22.82 0.0046 0.138 24.48 0.0052
SelfDeblur $0.732\pm 0.147$ $9.05\pm 2.46$ 0.1088 $0.733\pm 0.094$ $8.87\pm 2.21$ 0.0890
MPRNet $0.292\pm 0.101$ $\underline{22.42}\pm 3.21$ 0.0467 $0.334\pm 0.068$ $23.23\pm 2.19$ 0.0511
DeblurGANv2 $0.309\pm 0.111$ $\mathbf{22.55}\pm 3.44$ 0.0411 $0.325\pm 0.145$ $26.61\pm 3.28$ 0.0227
Pan-l0 $0.389\pm 0.090$ $14.10\pm 2.51$ 0.1961 $0.265\pm 0.082$ $20.68\pm 3.81$ 0.1012
Pan-DCP $0.325\pm 0.105$ $17.64\pm 3.63$ 0.1323 $0.235\pm 0.066$ $24.93\pm 3.60$ 0.0490
BlindDPS $0.246\pm 0.077$ $20.93\pm 2.09$ 0.0153 $0.216\pm 0.076$ $25.96\pm 2.45$ 0.0205
BIRD $0.294\pm 0.076$ $19.23\pm 1.88$ 0.0491 $0.212\pm 0.055$ $21.95\pm 1.62$ 0.0414
GibbsDDRM $0.293\pm 0.099$ $20.52\pm 2.81$ 0.0746 $0.216\pm 0.046$ $\underline{27.03}\pm 1.87$ 0.0430
LADiBI (Ours) $\underline{0.230}\pm 0.076$ $20.96\pm 2.34$ $\mathbf{0.0084}$ $\underline{0.197}\pm 0.071$ $21.08\pm 2.71$ $\mathbf{0.0068}$
GibbsDDRM* $\textbf{0.199}\pm 0.110$ $22.36\pm 3.79$ 0.0309 $\textbf{0.155}\pm 0.049$ $\textbf{27.65}\pm 2.66$ 0.0252
Table 6: Quantitative results on motion and Gaussian deblurring on AFHQ. Mean $\pm$ standard deviation shown.
Method Motion Gaussian
DPS w/ GT kernel 0.367 21.28 0.1120 0.330 23.52 0.0813
SelfDeblur $0.742\pm 0.158$ $9.04\pm 1.84$ 0.0650 $0.736\pm 0.112$ $8.84\pm 1.45$ 0.0352
MPRNet $0.324\pm 0.095$ $\underline{22.09}\pm 3.01$ 0.0382 $0.379\pm 0.081$ $21.97\pm 2.55$ 0.0461
DeblurGANv2 $0.323\pm 0.105$ $\mathbf{22.74}\pm 2.89$ 0.0350 $0.340\pm 0.084$ $\mathbf{27.12}\pm 2.94$ 0.0073
Pan-l0 $0.414\pm 0.133$ $14.16\pm 3.97$ 0.1590 $0.276\pm 0.079$ $21.04\pm 3.39$ 0.0320
Pan-DCP $0.371\pm 0.147$ $17.63\pm 5.94$ 0.1377 $0.297\pm 0.086$ $\underline{25.11}\pm 3.68$ 0.0263
BlindDPS $0.393\pm 0.061$ $20.14\pm 1.67$ 0.0913 $0.330\pm 0.057$ $24.79\pm 1.76$ 0.0268
BIRD $0.438\pm 0.110$ $18.92\pm 2.04$ 0.0286 $0.320\pm 0.079$ $21.87\pm 1.95$ 0.0089
GibbsDDRM $0.303\pm 0.114$ $19.44\pm 3.58$ 0.0265 $0.257\pm 0.121$ $24.01\pm 4.77$ $\mathbf{0.0040}$
LADiBI (Ours) $\mathbf{0.262}\pm 0.096$ $21.20\pm 2.72$ $\mathbf{0.0132}$ $\mathbf{0.204}\pm 0.066$ $24.33\pm 1.73$ 0.0065
GibbsDDRM* $\underline{0.278}\pm 0.099$ $19.00\pm 3.09$ 0.0180 $\underline{0.224}\pm 0.091$ $21.62\pm 3.53$ 0.0040
@@
Figure 11: Qualitative results on colorization task.
Figure 12: Qualitative results on non-linear deblurring task.
Figure 13: Demo inverse problem solving on car images.
Figure 14: Demo inverse problem solving on landscape images.
Figure 15: Qualitative samples on the RealBlur dataset.
@@ Appendix C Additional Ablation Study
Table 7: Quantitative comparison between linear and deep operator architecture
AFHQ Motion Blur
Method LPIPS $\downarrow$ PSNR $\uparrow$ KID $\downarrow$
Panl0 0.414 14.16 0.1590
PanDCP 0.371 17.63 0.1377
SelfDeblur 0.742 9.04 0.0650
MPRNet 0.324 22.09 0.0382
DeblurGANv2 0.323 22.74 0.0350
BlindDPS 0.393 20.14 0.0913
GibbsDDRM 0.303 19.44 0.0265
GibbsDDRM* 0.278 19.00 0.0180
LADiBI (Kernel Operator) 0.262 21.20 0.0132
LADiBI (U-Net Operator) 0.343 18.93 0.0146
@@
In this section, we provide additional ablation study on various hyperparameters and components of our LADiBI algorithm. All ablation studies are conducted on the motion deblurring task using the AFHQ dataset.
Figure 17 presents both the LPIPS and KID score for the value of the timestep $T_{s}$ at which we encode the measurement using SDEdit [Meng et al., 2022]. Our experiments align with the observation in Meng et al. [2022]: If $T_{s}$ is too large most of the information about the measurement has been replaced by noise which does not allow the sampling process to leverage useful features of the degraded image. On the other hand, if $T_{s}$ is too small, the sampling process is not equipped with enough scheduled steps to reach the target distribution.
@@
Finally, we evaluate the performance of our algorithm when using a unconstrained configuration for $^{\mathcal{A}}_{\hat{\phi}}$ in constrained tasks. In particular, we test the neural network architecture for the operator in linear inverse problems and we showcase quantitative results for the motion deblurring task in Table 7, and qualitative samples in Figure 18. We observe that, although performing worse than LADiBI employed with the aligned linear operator architecture, the neural operator is still capable of producing estimates of decent quality while also preserving a design that allows applicability to highly unconstrained image restoration task.
Appendix D Prior Coverage of Large Pre-trained Text-to-image Diffusion
In this section we provide an empirical justification for using Stable Diffusion v1.4 as our base model. To demonstrate this property, Figure 19 provides qualitative samples of images with Gaussian blur, motion blur and after JPEG compression. These examples show that Stable Diffusion 1.4 already capture rich distributions of both target images and common degradation artifacts.
(a) Generated images with Gaussian blur effect.
(b) Generated images with motion blur effect.
(c) Generated images with JPEG compression effect.
Figure 19: Drawn samples from the posterior distribution using our baseline model.
Appendix E Limitations and Future Works